# Emergence of Gibbs ensemble as the unique steady state in open quantum systems

### v1

@@ Appendix A More details for the Fredkin spin chain with boundary driving
Next we give an explanation of Eq.(10) in the main text, where the number of paired parenthesis is given by
$N^{k}=\sum_{i=1}^{N-1}\sum^{\lfloor\frac{N-1-i}{2}\rfloor}_{m=0}P^{\uparrow}_{i}P^{\downarrow}_{i+2m+1}\prod_{j=0}^{m-1}\left(1-P^{\uparrow}_{i}P^{\downarrow}_{i+2j+1}\right)\left(1-P^{\uparrow}_{i+2(m-j)}P^{\downarrow}_{i+2m+1}\right).$ (12)
In order to count paired parenthesis, we should make sure it is a pair of left and right parenthesis, so here we have $P^{\uparrow}_{i}P^{\downarrow}_{i+2m+1}$, which is separated by $2m$ distance. Moreover, they cannot pair with any parenthesis inside. Note that if they are separated by odd distance, then one of them must be paired with one parenthesis between them. By projecting out unwanted configurations using projectors in product form, we arrive at (12).
Now we discuss the full counting statistics (FCS) of the Fredkin model with boundary driving. Since $[\tilde{H},S^{z}]=0$, the FCS of operator $S^{z}=\sum_{i}\sigma^{z}_{i}$ is $\tilde{G}(\alpha)\propto\text{Tr}\left(e^{-\beta\tilde{H}+i\alpha S^{z}}\right)$ and the full counting function $G(\alpha)=\tilde{G}(\alpha)/\tilde{G}(0)$. From the form of $G$ we know $\text{Re}[G(\alpha)]=\text{Re}[G(-\alpha)]$ and $\text{Im}[G(\alpha)]=-\text{Im}[G(-\alpha)]$. Next we calculate the simple case $\text{Dis}=\{\sqrt{\gamma_{1}}L_{1},\sqrt{\gamma_{2}}L_{1}^{\dagger},\sqrt{\gamma_{2}}L_{N},\sqrt{\gamma_{1}}L_{N}^{\dagger}\}$ where $L_{i}=\sigma^{+}_{i}$. The size of a sector $\mathcal{K}(k,b,a)$ is (Salberger and Korepin, 2017) $\text{D}_{N}(\mathcal{K})=\left(\begin{array}[]{c}N\\ N-k\end{array}\right)-\left(\begin{array}[]{c}N\\ N-k+1\end{array}\right)$ and all possible $(k,b,a)$ that satisfies $2k+b+a=N$ appear exactly once. The FCS is
@@
$\tilde{H}=\frac{\beta_{l}}{\beta_{l}+\beta_{r}}N_{a}+\frac{\beta_{r}}{\beta_{l}+\beta_{r}}N_{b}-\frac{1}{2}\left(\frac{3\beta_{l}}{\beta_{l}+\beta_{r}}-1\right)(-1)^{N^{b}}-\frac{1}{2}\left(\frac{3\beta_{r}}{\beta_{l}+\beta_{r}}-1\right)(-1)^{N^{a}}\,,$ (14)
where $\beta_{l}\equiv\log(\frac{\gamma_{1}}{\gamma_{2}})$, $\beta_{r}\equiv\log(\frac{\gamma_{4}}{\gamma_{3}})$ and $\beta=\frac{\beta_{l}+\beta_{r}}{2}$ (Here we use a different notation for $\beta_{r}$, which is differed by a minus sign in comparison with the definition in the main text.), or equivalently $T=\frac{2T_{l}T_{r}}{T_{l}+T_{r}}$. When $\beta_{l}=\beta_{r}$ (14) is reduced to (11) differed by a constant factor. We can similarly calculate the FCS of magnetization
@@ B.1 Spin-1 $t-J_{z}$
$H_{t-J_{z}}=\sum_{i}\left(S^{+}(1-P^{\downarrow})\right)_{i}\left(S^{-}P^{\uparrow}\right)_{i+1}+\left(S^{+}P^{\downarrow}\right)_{i}\left(S^{-}(1-P^{\uparrow})\right)_{i+1}+h.c.\,.$ (16)
@@
$L_{\text{Dis2}}=\{L_{i,+}\equiv\frac{\sqrt{\gamma_{1}}}{2}\left(P^{\uparrow}_{i}S^{+}_{i+1}+S^{+}_{i}P^{\uparrow}_{i+1}\right),\,L_{i,-}\equiv\frac{\sqrt{\gamma_{2}}}{2}\left(P^{\downarrow}_{i}S^{-}_{i+1}+S^{-}_{i}P^{\downarrow}_{i+1}\right),\,\frac{\sqrt{\gamma_{2}}}{\sqrt{\gamma_{1}}}L_{i,+}^{\dagger},\,\frac{\sqrt{\gamma_{1}}}{\sqrt{\gamma_{2}}}L_{i,-}^{\dagger}\}$ (18)
@@
$L_{\text{Dis3}}=\{L_{0,+}\equiv\sqrt{\gamma_{1}}S_{1}^{+},L_{0,-}\equiv\sqrt{\gamma_{2}}S_{1}^{-},L_{i,+}\equiv\sqrt{\gamma_{3}}(S_{i}^{+})^{2}(S_{i}^{-})^{2},L_{i,-}\equiv\sqrt{\gamma_{3}}(S_{i}^{-})^{2}(S_{i}^{+})^{2}\}$ (19)
@@ B.2 Spin-1 pair-flip
The Hamiltonian of spin-1 pair-flip model reads (Moudgalya and Motrunich, 2022)
@@
$L_{\text{Dis}}=\{L_{i,+}\equiv\sqrt{\gamma_{1}}S_{2i+1}^{+},\,L_{i,-}\equiv\sqrt{\gamma_{1}}S_{2i+2}^{-},\,\frac{\sqrt{\gamma_{2}}}{\sqrt{\gamma_{1}}}L_{i,+}^{\dagger},\,\frac{\sqrt{\gamma_{2}}}{\sqrt{\gamma_{1}}}L_{i,-}^{\dagger}\}\,,$ (22)
@@ B.3 Spin-1 dipole conserving
The Hamiltonian of spin-1 dipole conserving model (Sala et al., 2020) reads
@@
$L_{\text{Dis}}=\{L_{0,+}\equiv\sqrt{\gamma_{1}}S_{1}^{+},L_{0,-}\equiv\sqrt{\gamma_{2}}S_{1}^{-},L_{i,+}\equiv\sqrt{\gamma_{3}}S_{i}^{-}S_{i+1}^{+}P_{i,i+1},L_{i,+}\equiv\sqrt{\gamma_{4}}S_{i}^{+}S_{i+1}^{-}P_{i,i+1}\}$ (25)

### v2

@@ Appendix A More details for the Fredkin spin chain with boundary driving
Next we give an explanation of Eq.(10) in the main text. In order to count paired parenthesis, we should make sure it is a pair of left and right parenthesis, so here we have $P^{\uparrow}_{i}P^{\downarrow}_{i+2m+1}$, which is separated by $2m$ distance. Moreover, they cannot pair with any parenthesis inside. Note that if they are separated by odd distance, then one of them must be paired with one parenthesis between them. By projecting out unwanted configurations using projectors in product form, we arrive at Eq.(10).
Now we discuss the full counting statistics (FCS) of the Fredkin model with boundary driving. Since $[\tilde{H},S^{z}]=0$, the FCS of operator $S^{z}=\sum_{i}\sigma^{z}_{i}$ is $\tilde{G}(\alpha)\propto\text{Tr}\left(e^{-\beta\tilde{H}+i\alpha S^{z}}\right)$ and the full counting function $G(\alpha)=\tilde{G}(\alpha)/\tilde{G}(0)$. From the form of $G$ we know $\text{Re}[G(\alpha)]=\text{Re}[G(-\alpha)]$ and $\text{Im}[G(\alpha)]=-\text{Im}[G(-\alpha)]$. Next we calculate the simple case $\text{Dis}=\{\sqrt{\gamma_{1}}L_{1},\sqrt{\gamma_{2}}L_{1}^{\dagger},\sqrt{\gamma_{2}}L_{N},\sqrt{\gamma_{1}}L_{N}^{\dagger}\}$ where $L_{i}=\sigma^{+}_{i}$. The size of a sector $\mathcal{K}(k,b,a)$ is [49] $\text{D}_{N}(\mathcal{K})=\left(\begin{array}[]{c}N\\ N-k\end{array}\right)-\left(\begin{array}[]{c}N\\ N-k+1\end{array}\right)$ and all possible $(k,b,a)$ that satisfies $2k+b+a=N$ appear exactly once. The FCS is
@@
$\tilde{H}$ $=\frac{\beta_{l}}{\beta_{l}+\beta_{r}}N_{a}+\frac{\beta_{r}}{\beta_{l}+\beta_{r}}N_{b}-\frac{1}{2}\left(\frac{3\beta_{l}}{\beta_{l}+\beta_{r}}-1\right)(-1)^{N^{b}}$ (13)
$-\frac{1}{2}\left(\frac{3\beta_{r}}{\beta_{l}+\beta_{r}}-1\right)(-1)^{N^{a}}\,,$
where $\beta_{l}\equiv\log(\frac{\gamma_{1}}{\gamma_{2}})$, $\beta_{r}\equiv\log(\frac{\gamma_{4}}{\gamma_{3}})$ and $\beta=\frac{\beta_{l}+\beta_{r}}{2}$ (Here we use a different notation for $\beta_{r}$, which is differed by a minus sign in comparison with the definition in the main text.), or equivalently $T=\frac{2T_{l}T_{r}}{T_{l}+T_{r}}$. When $\beta_{l}=\beta_{r}$ (13) is reduced to Eq.(11) differed by a constant factor in the main text. We can similarly calculate the FCS of magnetization
@@ B.1 Spin-1 $t-J_{z}$
$H_{t-J_{z}}=$ $\sum_{i}\left(S^{+}(1-P^{\downarrow})\right)_{i}\left(S^{-}P^{\uparrow}\right)_{i+1}$ (15)
$+\left(S^{+}P^{\downarrow}\right)_{i}\left(S^{-}(1-P^{\uparrow})\right)_{i+1}+h.c.\,.$
@@
[…]
$L_{i,-}\equiv\frac{\sqrt{\gamma_{2}}}{2}\left(P^{\downarrow}_{i}S^{-}_{i+1}+S^{-}_{i}P^{\downarrow}_{i+1}\right),\,$
$\frac{\sqrt{\gamma_{2}}}{\sqrt{\gamma_{1}}}L_{i,+}^{\dagger},\,\frac{\sqrt{\gamma_{1}}}{\sqrt{\gamma_{2}}}L_{i,-}^{\dagger}\}$
@@
[…]
$L_{i,+}\equiv\sqrt{\gamma_{3}}(S_{i}^{+})^{2}(S_{i}^{-})^{2},$
$L_{i,-}\equiv\sqrt{\gamma_{3}}(S_{i}^{-})^{2}(S_{i}^{+})^{2}\}$
@@ B.2 Spin-1 pair-flip
The Hamiltonian of spin-1 pair-flip model reads [38]
@@
[…]
$\frac{\sqrt{\gamma_{2}}}{\sqrt{\gamma_{1}}}L_{i,+}^{\dagger},\,\frac{\sqrt{\gamma_{2}}}{\sqrt{\gamma_{1}}}L_{i,-}^{\dagger}\}\,,$
@@ B.3 Spin-1 dipole conserving
The Hamiltonian of spin-1 dipole conserving model [37] reads
@@
[…]
$L_{i,+}\equiv\sqrt{\gamma_{3}}S_{i}^{-}S_{i+1}^{+}P_{i,i+1},$
$L_{i,+}\equiv\sqrt{\gamma_{4}}S_{i}^{+}S_{i+1}^{-}P_{i,i+1}\}$